3. Suppose that A and B are two events defined over the same sample space, with probabilities P(A) 3/4 and P(B)- 3/8. (a) Show that P(A UB) 2 3/4. (b) Show that 1/8 < P(AB) 3/8 (c) Give inequalities analogous to (a) and (b) for P(A) 2/3 and P(B)1/2.
Answer #1

Given the probabilities of 2 events, P (A)-3/4. P (B)-3/8 .

Use the identities P(AB)= P (B) P (AIB) and P(AUB)=P(A) P (B)-P(AB)

a) Now rewriting the second identity,

P(AUB) P(A) P (B) - P (B) P (AB) P (A U B) = P (A) P (B) [1-P (AB) P (A U B) P (A) since 0 < P (AB) 1

The proof is complete.

b) Again using the identities,

P (A) P (B) P(AB) 3/4 3/4+1/8- P (AB) 3/4 1 /8 P (AB)

P (A n B) = P (AB) P (B) P (A n B) < P (B) since 0 < P (AIB) P(AnB) 3/8 1

Thus {color{Blue} 1/8 leqslant Pleft ( Acap B ight ) leqslant 3/8}

c) When P (A) =2/3. P (B)-1/2 . Here P (A) = 2/ 3 is larger, so the analogous inequalities are

{color{Blue} Pleft ( Acup B ight ) geqslant 2/3}

Pleft ( Acup B ight )leqslant 2/3 Pleft ( A ight )+Pleft ( B ight ) -Pleft ( AB ight ) leqslant2/3 2/3+1/2-Pleft ( AB ight ) leqslant 2/3 {color{Blue} 1/2leqslant Pleft ( AB ight )}

P (A n B) = P (AB) P (B) P (An B) < P (B) since P (A|B) < 0

Thus {color{Blue} 1/2 leqslant Pleft ( Acap B ight ) leqslant 2/3}